OpenStudy (zmudz):

Let F(x) be the real-valued function defined for all real x except for x = 0 and x = 1 and satisfying the functional equation \(F(x) + F((x-1)/x) = 1+x.\)Find the F(x) satisfying these conditions. Write F(x) as a rational function with expanded polynomials in the numerator and denominator. Help needed! Thank you!

ganeshie8 (ganeshie8):

\[F(x)+F\left(\frac{x-1}{x}\right) =1+x\tag{1}\] replace \(x\) by \((x-1)/x\) in \((1)\) and get \[F\left(\frac{x-1}{x}\right)+F\left(\frac{-1}{x-1}\right) =1+\frac{x-1}{x}\tag{2}\] replace \(x\) by \(-1/(x-1)\) in \((1)\) and get \[F\left(\frac{-1}{x-1}\right)+F\left(x\right) =1+\frac{-1}{x-1}\tag{3}\]

ganeshie8 (ganeshie8):

If you let \(F(x)=a\), \(F\left(\frac{x-1}{x}\right)=b\) and \(F\left(\frac{-1}{x-1}\right)=c\), the previous equations become : \[a+b = 1+x\tag{1'}\] \[b+c = 1+\frac{x-1}{x}\tag{2'}\] \[c+a = 1+\frac{-1}{x-1}\tag{3'}\]

ganeshie8 (ganeshie8):

3 equations and 3 unknowns (a,b,c) you can solve them